what is square root of 13 over 28?
Answer by edjones(8007)

sqrt(13/28)
=sqrt(13)/sqrt(7*4)
=sqrt(13)/2sqrt(7)
=sqrt(13)/2sqrt(7) * sqrt(7)/sqrt(7) so an irrational number wont be in the denominator.
=sqrt(91)/14
.
Ed